User name jpchen
Log entry time 09:44:22 on March 5,2007
Entry number 194825
keyword=sieve out field to 0%: 2.635
Continue and completed optics study (0% raster on, then raster off,
moved out sieve slit on left arm, changed field to 0%: 2.635 GeV/c
(did cycle of Q2/Q3).
Ready to start current scan on 12C.